### v1

# Conceptualization, Methodology, Writing, Visualization

Abstract
With the large-scale integration of electric vehicles (EVs) in the distribution grid, the unpredictable nature of EV charging introduces considerable uncertainties to the grid’s real-time operations. This can exacerbate load fluctuations, compromise power quality, and pose risks to the grid’s stability and security. However, due to their dual role as controllable loads and energy storage devices, EVs have the potential to mitigate these fluctuations, balance the variability of renewable energy sources, and provide ancillary services that support grid stability. By leveraging the bidirectional flow of information and energy in smart grids, the adverse effects of EV charging can be minimized and even converted into beneficial outcomes through effective real-time management strategies. This paper explores the negative impacts of EV charging on the distribution system’s real-time operations and outlines methods to transform these challenges into positive contributions. Additionally, it provides an in-depth analysis of the real-time management system for EV charging, focusing on state estimation and management strategies.
keywords:
[…]
1 Introduction
As industrial society continues to evolve, the transportation sector has remained a major contributor to greenhouse gas (GHG) emissions fachrizal2021combined . While traditional internal combustion engine vehicles have facilitated convenient mobility, they have also significantly polluted the environment. In recent years, as the transportation sector strives for cleaner and more sustainable solutions, electric vehicles (EVs) have become a key factor in reducing GHG emissions. With their low maintenance requirements and superior performance, the number of EVs has increased rapidly, marking significant progress in the electrification of transportation zenhom2024comprehensive . According to the International Energy Agency’s Global EV Outlook 2024, the global stock of EVs, excluding two- and three-wheelers, is projected to grow from under 45 million in 2023 to 250 million by 2030, and further to 525 million by 2035. By that time, more than a quarter of all vehicles on the road will be electric. The widespread adoption of EVs offers numerous benefits, such as reducing GHG emissions through decreased fossil fuel use and enhancing transportation capacity to lower costs. However, integrating EVs on a large scale into distribution networks poses significant challenges for grid operation. It is anticipated that by 2040, EVs will account for approximately 28% of the market share, resulting in an 11-20% increase in global electricity consumption kapustin2020long . Additionally, the increased demand from EV charging during peak hours could place greater pressure on the safe and stable operation of existing distribution system. This could also result in higher cost, as consumers are often required to pay substantial charges based on time-of-use tariffs engel2022hierarchical . To address these challenges, it is crucial to actively integrate EVs into energy infrastructure and management systems. With their dual characteristics serving as controllable loads or energy storage devices li2021coordinating , EVs can help smooth load fluctuations in the power grid, balance the intermittency of distributed generation (DG), and provide ancillary services to maintain grid stability. Therefore, through appropriate real-time management strategies, the negative impacts of EV charging can be minimized, or even transformed into positive outcomes. In this regard, this paper first illustrates the negative impacts of EV charging on the real-time operation of distribution system. It then focuses on mitigating these impacts from two perspectives:
  • 1)
    Providing a comprehensive summary of management strategies to regulate EV impacts, such as smart charging, charging environment management, energy coordination, battery management, and ancillary services.
  • 2)
    Conducting an in-depth analysis of the real-time management system for EV charging by focusing on real-time state estimation of the distribution network and the management of EV charging activities.
The overall structure of this paper is depicted in Figure 1. The rest of this paper is organized as follows. Section 2 discusses the negative impacts of EV charging on distribution systems. Sections 3 and 4 examine the management strategies and the real-time management system for EV charging to mitigate these impacts, respectively. Finally, Section 5 offers conclusions and insights for future research.
Figure 1: Summary of research on EV charging’s impacts reviewed in this study.
2 Negative impacts of EV charging on distribution systems
2.1 Power quality
2.1.1 Voltage
Based on the characteristics of the power system, voltage is closely related to reactive power, which is more difficult to generate and inject into grids than active power sayed2022electric . Thus, the high demand for reactive power from EV nonlinear charging can significantly impact the real-time voltage of grid operation. The direct impact of large-scale EV charging on distribution system operation is a voltage drop at the coupling point, which may lead to voltage deviation exceeding the regulatory requirements. For instance, rahman2022comprehensive investigates the factors affecting voltage distribution, such as locations of power sources and EV penetration levels, and compares conditions of multiple parallel load lines with unequal loads. Furthermore, continuous violations may lead to grid operation instability and could even result in blackouts. Therefore, when voltage violations or frequency violations exceed specified limits, it’s necessary to adopt corrective measures to restore violations to normal levels in order to avoid damage to power equipment and negative effects on grid operation safety.
Regarding voltage regulation, reactive power dispatch and load demand management are effective methods to control voltage drops wu2012pev ; leemput2015reactive ; sousa2015multi . Some studies wang2012load ; li2012impacts compare the effects of random charging and smart charging on the voltage variation of the distribution system, finding that smart charging can increase EV penetration rates. Moreover, another impact on voltage is the issue of three-phase voltage unbalance, which is primarily caused by single-phase charging. Uneven distribution of charging loads across three phases may lead to a severe condition of three-phase voltage unbalance yong2015review . richardson2010impact performs a detailed study on this issue by connecting all EVs to a single phase and confirms the seriousness of the problem. To address this problem, a proper load management strategy is usually adopted to mitigate three-phase voltage unbalance by evenly dispatching EV charging loads across three phases. Figure 2 presents the process of voltage deviation from EV charging and regulating voltage to the normal level kersting2018distribution . In Figure 2, $V_{s}$, $V_{load}$ and $V_{manage}$ represent the initial voltage of the node, the voltage after applying EV charging and loads, and the voltage after exerting management methods, respectively. $RI$ and $jXI$ represent the real and imaginary parts of the impedance of EV charging and loads, respectively. $RI_{cap}$ and $jXI_{cap}$ represent the real and imaginary parts of the impedance of the reactive power compensation supply, respectively.
Figure 2: The phasor diagram for regulating voltage deviation to the normal level.
2.1.2 Harmonics
Power electronic devices, essential for EV charging, are extensively utilized in this field. However, their use can lead to power quality issues in the distribution system, such as the introduction of harmonics, which affect real-time operations.
As the foundation of EV charging, power electronic devices have been widely used. This has caused power quality degradation in the real-time operation of the distribution system, such as harmonics yong2015review . nguyen2013harmonic shows similar outcomes: EV charging injects significant harmonics into the distribution system and can cause unacceptable Total Harmonic Distortion (THD) of voltage. Regarding studies on the effect of harmonics, establishing models of EV chargers is a critical point, where gatta2016pq develops a sensitivity analysis about the composition of harmonic disturbances due to AC/DC converters installed in EVs, and rahman2022comprehensive compares simulation results for different power converter topologies in steady-state operation.
The basic configuration of an EV charger uses a back-to-back converter structure. On the input side, a diode bridge rectifier determines the current, which is highly peaked and dominated by low-order harmonics. This leads to voltage deviations, reduced power quality, and de-rating of system components yong2015review ; rahman2022comprehensive . The peak current is superimposed on the sinusoidal current drawn by the EV charger and other loads in the distribution system. It produces a non-sinusoidal voltage drop across the grid impedance. Thus, both the coupling point and the distribution grid contain additional harmonics. The effect of the above condition depends on the parameters of the distribution line; if the grid impedance is small, the voltage drop at the coupling point due to non-sinusoidal current is small. Although EV charging may bring harmonic pollution to the power grid, employing filtering and advanced power electronic devices can alleviate this problem. For instance, rai2019bridgeless proposes a single-ended primary inductor converter for power factor correction operation, and rahman2022comprehensive adopts a boost power factor correction circuit along with the diode bridge rectifier to solve the above condition, through improving and regulating the rectified voltage to generate a minimal ripple DC voltage.
2.2 Stable operation of distribution system
Large-scale uncoordinated charging of EVs may negatively affect the real-time operation of the distribution system, causing overloading, voltage drops, power outages, and posing a threat to the stable and safe operation of the distribution network hasanien2015adaptive . This stable problem can be written as:
For any initial state $x(t_{0})$ with $\left\|x\left(t_{0}\right)\right\|<k_{1}$ and any input $u(t_{\mathrm{s}})$ with $\sup_{t_{\mathrm{s}}\geq t_{0}}\|u(t_{\mathrm{s}})\|<k_{2}$, $x(t_{\mathrm{s}})$ exists and the output $y{(t_{s})}$ satisfies
$\begin{gathered}\|y(t_{\mathrm{s}})\|\leq\beta\left(\left\|x\left(t_{0}\right)\right\|,t_{\mathrm{s}}-t_{0}\right)+\alpha\left(\sup_{t_{0}\leq\tau\leq t_{\mathrm{s}}}\|u(\tau)\|\right),\end{gathered}$ (2)
where all $t_{\mathrm{s}}\geq t_{0}\geq 0$.
Research on the impact of EV charging on the stability of the distribution system typically focuses on three aspects: rotor angle stability, frequency stability, and voltage stability. References dharmakeerthi2014impact ; shi2012dynamic ; zhou2016assessment indicate that EV charging reduces the level of power system stability, while references wu2011transient ; el2009advantages suggest that the Vehicle-to-Grid (V2G) model can improve the system stability level.
2.2.1 Peak loads
Both electricity consumption and EV charging loads are closely related to human activities. Without controlled charging, EV charging demand may overlap with existing peak loads in the grid fachrizal2020smart . This overlap would further intensify grid loads during peak time, thereby stressing grid operation and posing risks to the security and stability of the distribution system. Peak load increase generated by EV charging has become a critical factor for grid operation and risk assessment stiasny2021sensitivity . Some studies have focused on this issue. For instance, weiller2011plug suggests that, during commuting hours, new load peaks could exceed natural peaks if EV charging loads are not sufficiently shifted to off-peak periods. Another study shafiq2020reliability indicates that uncontrolled EV charging, especially during peak time, could lead to up to 6.89% load loss. Since traditional distribution systems are designed to handle peak loads, reducing peak demand can also significantly lower overall construction costs. Key factors affecting peak loads in the distribution system include EV charging time, charging location, charging power, and penetration rates of EV charging. Thus, the load management strategy can help balance power loads and reduce the difference between peak and valley loads fachrizal2020smart . Common load management methods include off-peak and valley-filling charging, which shifts EV charging from peak to lower-demand times. This charging approach avoids charging during peak periods and fills low consumption periods, reducing system loss and improving load factor sortomme2010coordinated .
Regarding the impact of EV-grid-connected charging on peak loads in practical applications, many countries, including the United States and Germany, have analyzed the effects of EV charging on load distribution based on their specific circumstances and have proposed corresponding solutions. These solutions include delayed charging methods weiller2011plug , using EVs as stable power storage devices in the grid hartmann2011impact , shifting EV charging to nighttime hours mullan2011modelling , transferring EV loads from peak to off-peak periods by implementing demand response (DR) strategies park2013impact and utilizing V2G reverse power flow to reduce peak load in the grid. Specifically, mosaddegh2017optimal developed a Distribution Optimal Power Flow (DOPF) model incorporating a neural network model of controllable loads to mitigate peak loads. Based on a microgrid energy management system framework, solanki2015including proposes optimal dispatch strategies for dispatchable generators, energy storage systems, and controllable peak loads to achieve peak load dispatch effectively.
2.2.2 The coping capacity of EV charging areas to grid faults
Faults in grids may cause fast variations of voltage and bursts of harmonics, as well as a serious influence on EV charging. For instance, faults causing undervoltage at or below 0.3 p.u. can lead to ceasing charging for 2–10 seconds mckillop2023impact . onar2010grid compares conditions of a three phase-ground fault and after fault clearance, obtaining the conclusion that systems connected to EVs are more sensitive to disturbances and less stable in magnitude deviations and adjustment time. mckillop2023impact investigates the impact of EVs on power grids during network faults through testing EVs’ responses to a double-line-to-ground transmission fault, obtaining the result of successive over-frequency and under-frequency of 50.78 and 49.22 Hz, respectively. Therefore, regarding fault conditions in grids, the ability of EV charging areas to ride through grid faults and mitigate grid faults is worth attention. EV charging areas with low-voltage ride-through (LVRT) function can effectively handle faults and prevent system instability onar2010grid , and V2G mode can also be used for grid support during faults according to LVRT requirements katic2019impact . For instance, katic2019impact tests six characteristic types of faults to demonstrate the positive impact of LVRT. Table 1 summarizes the research on enhancing EV charging areas’ coping capacity to grid faults.
Table 1: Summary of the research on enhancing EV charging areas’ coping capacity to grid faults
Ref. no. Year Technical foundation Implementation approach Aims and resolved issues
mishra2022intelligent 2022 Least mean square algorithm A reconfigurable multi-objective charging control architecture Overcome various grid abnormalities during EV charging operation
baghaee2019anti 2020 Support vector machines Anti-islanding protection scheme for low voltage-sourced converter-based microgrids Islanding and grid-fault detection.
shin2012development 2012 Wireless sensor network Smart grid monitoring system connected with EV charging system using anti-islanding method EV charging process continues without any serious fault
onar2010grid 2018 Dynamic combination of EV chargers and single-phase induction motors Implement an LVRT scheme to inject reactive power into the grid and regulate EV charging rate Handle faults and prevent dynamic voltage instability.
almeida2015electric 2015 Combination of inertial emulation and droop control Primary frequency control technique with EVs Safe integration of intermittent renewable energy sources
pandit2024frequency 2024 MM-SFR model Propose a non-linear optimization framework incorporating constraints of EV aggregator, frequency security, converter voltage, security, and LVRT constraints Develop a framework for quantifying EVs’ contribution to providing frequency support
mckillop2023impact 2023 Fault condition test Test EVs physically under various network fault conditions using a grid simulator supply interface Characterize fault ride-through (FRT) performances
katic2019impact 2019 V2G mode according to the LVRT requirements Test the possibility of voltage improvement during voltage dips in V2G mode Provide grid support during faults through V2G mode
yadav2023low 2023 LVRT Applications of LVRT Boost the resilience of the power network against extreme events
falahi2013potential 2013 V2G services Pair up a photovoltaic source with an EV charger through a single-phase bidirectional charger topology Study the potential of EVs to help PV sources during LVRT
dietmannsberger2017simultaneous 2017 A three-phase inverter model Inverters with both LVRT capability and anti-islanding protection simultaneously Solve the conflict between LVRT capability and anti-islanding detection requirements
3 Management methods of regulating impacts from EV charging
3.1 Smart charging
In response to the growing number of EVs and their impacts on infrastructure, it is necessary to implement intelligent and coordinated charging management methods engel2022hierarchical . Many studies have focused on optimizing charging management through smart charging approaches, typically integrated with an energy management system (EMS), which is designed to determine optimal charging schemes and regulation strategies that lead to positive effects on the distribution network. Figure 3 presents the negative impacts from EV charging on net loads, and the positive impacts from smart charging. Through establishing models for controllable loads solanki2015including ; hafez2016integrating ; mosaddegh2017optimal , these models are integrated into the distribution system operation framework or EMS solanki2017sustainable ; wi2013electric ; van2016energy ; solanki2015including to determine the optimal charging strategies and dispatch decisions wi2013electric ; van2016energy ; engel2022hierarchical ; solanki2015including ; hafez2016integrating ; kiviluoma2011methodology ; zheng2018online . The advantage of this approach is optimizing EV charging management together with other components of power system. Determining the optimal charging strategies and optimal dispatch decisions is typically formulated as an optimization problem considering multiple factors, with the charging management objective encoded as a cost function engel2022hierarchical . Specifically, in solanki2015including , an EMS framework is proposed to determine optimal scheduling decisions considering dispatchable generators, energy storage systems, and peak demand for controllable loads. zheng2018online transforms the optimal charging problem to an optimal power flow problem to minimize the total system energy cost, then utilizes a modified convex relaxation technique to obtain the globally optimal solution. The key issue in establishing EV load models is reducing the uncertainties caused by EV charging behaviors. Due to the complexity of controllable loads and the limited data, it is difficult to model these loads using basic physical laws mosaddegh2017optimal . Common methods include probabilistic approaches papadopoulos2010predicting ; ramadhani2020review , stochastic optimization van2016energy ; liao2015dispatch , and evolutionary algorithms, such as neural networks hafez2016integrating ; mosaddegh2017optimal , particle swarm optimization regulski2014estimation , or genetic algorithms ju1996nonlinear . Moreover, based on these technical foundations, many studies have focused on developing efficient and accurate real-time management models to address the impact of EV integration on the real-time operation of distribution networks. A detailed introduction to this topic is provided in Section 4.
Figure 3: (a) The effects of EV charging on the electricity load profile on a day in residential distribution grids. (b) The positive impacts of managing EV charging on net loads through PV coordination and smart charging methods. Figure are developed with data from berg2022data .
3.2 Charging environment management
In addition to charging stations, common EV charging environments include workplaces and households, where rooftop photovoltaic (PV) systems are typically integrated with energy storage systems (ESS). Regarding workplace conditions, the net-zero energy building (NZEB) is a popular topic, and the synergy between EV loads and PV generation in buildings can help achieve this goal fachrizal2020smart . In the context of building energy management system with EV charging stations, engel2022hierarchical proposes a hierarchical economic model predictive control scheme for EV charging management, considering the objectives of monetary costs, building temperature comfort, EV charge satisfaction, and battery degradation. van2016energy designs a workplace energy management system with photovoltaic generation prediction and power flow optimization between PV systems, grids, and battery electric vehicles. lakshminarayanan2018real proposes a real-time optimal EM controller for V2G integration to provide an optimal schedule for the operation of the workplace microgrid system.
The issue of EV charging in households is closely related to the home energy management system (HEMS), which is necessary for residential electricity consumers to participate in DR programs actively ghazvini2017demand . shao2012grid proposes a DR strategy, and ghazvini2017demand presents an intelligent HEMS algorithm to schedule controllable appliances, i.e., EVs in smart homes. In addition, wi2013electric considers user preferences for determining EV charging schedules in smart homes, and shao2012grid introduces consumer comfort indices to measure the impact of demand response on consumers’ lifestyles.
3.3 Energy coordination and battery management
Renewable energy generation is flexible, environmentally friendly, and cost-effective, significantly reducing greenhouse gas emissions and environmental pollution. It is produced naturally and is subject to natural laws, which means it is inherently random and discontinuous. For example, wind power generation depends on variations in wind speed and direction, while photovoltaic generation is influenced by solar irradiance and shadow patterns, which are affected by geographical location, micro-climates, and seasons. Consequently, renewable energy generation is intermittent and fluctuating, posing challenges for power systems to maintain a real-time balance between supply and demand, thus requiring effective management. Additionally, since both electricity consumption and EV charging loads are closely linked to human activities, peak power consumption and EV charging demands are likely to occur simultaneously fachrizal2020smart . Distributed generation (DG) is often unstable and difficult to predict accurately. Coordinated EV charging can help balance power load demands at various times, thereby mitigating the intermittent and unstable effects of DG ghazvini2017demand ; wi2013electric ; van2016energy ; solanki2015including ; shepero2018modeling . This problem can be written as ramadhani2020review :
$\min$ $\sum_{t=t_{\mathrm{arr}}}^{t_{\mathrm{dep}}}\left(p_{t}+l_{t}-s_{t}-\mu_{\mathrm{tpark}}\right)^{2},$ (3)
s.t. $\eta_{x}\sum_{t=t_{\mathrm{arr}}}^{t_{\mathrm{dep}}}p_{t}\cdot\Delta t=\mathrm{SoC}_{\mathrm{tar}}-\mathrm{SoC}_{\mathrm{arr}},$ (4)
$0\leq p_{t}\leq p_{\max},$ (5)
where $t_{\text{arr }}$ and $t_{\text{dep }}$ are the arrival and departure times of EV respectively, $p_{t}$ is the charging power rate at time $t$, $l_{t}$ is the household load at time $t$, $s_{t}$ is solar power generation rate at time $t$, $\mu_{\mathrm{tpark}}$ is the mean net-load. In the constraint, $\eta_{p}$ is the charging efficiency, $\Delta t$ is the time step, $\mathrm{SoC}_{\mathrm{tar}}$ is the targeted state of battery, $\mathrm{SoC}_{\mathrm{arr}}$ is the state of battery at arrival time and $p_{\max}$ is the maximum charging power rate.
Given the high cost of energy storage systems, enhancing the utilization of DG through the control of flexible loads represents one of the effective solutions. An electric vehicle represents a movable battery storage load sayed2022electric , which can be used as a flexible and mobile storage unit. EV battery management has become a prominent topic, primarily focusing on battery modeling considering degradation engel2022hierarchical ; leippi2022review , and energy scheduling lakshminarayanan2018real ; acha2010effects ; liao2015dispatch ; sun2017optimal . With the large-scale deployment of EVs, their aggregated storage capacity can become a significant resource, no longer negligible acha2010effects .
3.4 Ancillary services
Due to the characteristics of electric vehicles, the EV battery can be considered both a power load and a distributed energy storage unit. Leveraging the energy storage function of the EV battery, EVs can provide power storage, power supply, and ancillary services to the distribution system when parked. With the modernization of power systems, integrating power and communication infrastructure based on smart grids enables energy exchange and information flow between EVs and the grid, which provides support for various services to enhance the reliability and sustainability of the power system. As a result, the V2G concept has emerged.
The ancillary services contain frequency regulation, spinning reserve, active power support, and reactive power compensation tan2016integration . Refs. rahman2022comprehensive ; tan2016integration ; guille2009conceptual ; kisacikoglu2010effects ; rotering2010optimal ; sortomme2011optimal ; ehsani2012vehicle pecifically discuss the implementation methods of V2G ancillary services. Ancillary services have the positive impact of maintaining the balance between the generation and load demand to ensure a stable and reliable power system. In addition, V2G services also help reduce emissions, increase profits, and provide additional income for electric vehicle owners. Meanwhile, using EVs to provide ancillary services presents several challenges. Beyond the issue of battery degradation from frequent bidirectional V2G operations, there is also the need for additional investment in bidirectional chargers.
3.4.1 Frequency regulation and Spinning reserve
Frequency regulation involves two modes: regulation up and regulation down pandit2024frequency ; almeida2015electric . These modes adjust the power levels of specific resources to fine-tune frequency by balancing supply and demand within a minute or less kempton2005vehicle . Specifically, when loads increase, EV charging rates can be reduced to achieve the regulation down mode in response to the increased load.
Spinning reserve provides synchronized generation capacity with fast response in a typical duration of minutes, generally within 10 minutes to compensate for the generation outage han2014practical ; bandyopadhyay2011aggregator . More specifically, EVs’ charging rates can be decreased by unidirectional V2G technologies to attain the additional spinning reserve.
3.4.2 Active power support and reactive power compensation
Active power support is a service to optimize load curves in the distribution system, which flattens the peak load profile by “peak load shaving” and “load leveling”. Based on bidirectional V2G technologies, excessive EV battery energy is discharged to shave off the peak load and alleviate the applied stress on the distribution system. Furthermore, this service also reduces overall power losses and additional equipment upgrade costs by shaving peak load and maintaining a lower power level in the distribution system birnie2009solar ; rotering2010optimal ; zhang2012integration .
Reactive power compensation is a technique to provide voltage regulation and power factor correction tan2016integration , further increasing power system operating efficiency and reducing power loss, such as guille2009conceptual ; kisacikoglu2010effects . In conventional reactive power supply methods, a capacitive reactive power is needed for this function. Distribution generators and static Volt–Ampere Reactive (VAR) compensators are most commonly used. Based on the capacitive reactive power reserved in the DC-link capacitor of the EV bidirectional battery charger, the EV can supply reactive power compensation by controlling the AC/DC converter without any battery degradation. Methods of managing EV charging for mitigating negative impacts on the distribution system are presented in Figure 4.
Figure 4: Methods of managing EV charging for mitigating negative impacts on the distribution system.
4 Real-time management system of EV charging
4.1 State estimation
The power system is a complex, large-scale, cyber-physical-social system where numerous components interact mutually to maintain a dynamic balance. Integration of EVs will disrupt the current balance, necessitating the establishment of a new equilibrium. As EVs become more prevalent, various parts of the distribution networks will react differently, leading to diverse outcomes and potential impacts. For instance, sayed2022electric tests different locations of connecting EVs and demonstrates that EV loads may cause line trips, cascading failures, multiple lines’ overloading, and even a blackout. Based on the topology of distribution networks, real-time state estimation can be conducted to assess the sensitivity of holding EV charging at different locations, and then weaknesses in the network can be identified. It is significant to further enhance monitoring and management of weaknesses during real-time operation to ensure a secure and stable distribution system. The general model of state estimation is written as:
$\begin{gathered}Z=h(X,Y)+N,\end{gathered}$ (7)
where $Z$ is the measurement vector of measurement, $X$ and $Y$ are vectors of state variables, $N$ is measurement noise, and $h$ is a function that relates state variables to measurements. Knowing the network topology is a prerequisite for evaluating the impact of EV charging on distribution networks taheri2019milp . However, due to the limited presence of meters and infrequently-calibrated line parameters, distribution grid operators have problems with accurate and updated real-time information on the grid topology, which causes the need for dynamic identification of network topology park2018exact . Studies alturki2018marginal ; taheri2019milp ; park2018exact have contributed to estimating grid topology according to data measured at smart meters and grid sensors sayed2022electric .
With knowledge of power grid topology, methods of analyzing grid stability, such as PV and QV curves, can be utilized to locate buses or nodes where the impact of EV charging can be amplified, thus achieving the identification of weaknesses in the network. Voltage stability indices can be utilized to solve this issue, which have two main classifications modarresi2016comprehensive : 1) Jacobian matrix and system variables based VSIs. 2) Bus, line, and overall VSIs. The voltage stability index for the bus is given by
$\begin{gathered}VSI_{i}=\left[1+\left(\frac{I_{i}}{V_{i}}\right)\left(\frac{\Delta V_{i}}{\Delta I_{i}}\right)\right]^{\alpha},\end{gathered}$ (9)
where $I_{i}$ and $V_{i}$ are the current and voltage at bus $i$ respectively, $\Delta I_{i}$ and $\Delta V_{i}$ are current and voltage deviation at bus $i$ respectively, and $\alpha$ is a constant number equal or greater than 1. For instance, the significance of EV charging location is presented by considering different feeders and load curves in rahman2022comprehensive . Specifically, alturki2018marginal determines the more impactful nodes in assessing network capacity by identifying the maximum possible change in each specific node. sayed2022electric simulates the impacts from increased loads on different buses, and obtains the result that buses 6 and 7 are less sensitive with smaller variations than other buses. deb2018impact tests the different buses’ holding capacity of EV charging by six different cases of EV charging placement and proposes a novel voltage index. It is demonstrated that the system has the capability to support the placement of fast charging stations at strong buses within a certain threshold. However, installing fast charging stations at weak buses adversely affects the stable operation of the power system. rahman2022comprehensive compares and analyzes the impact of EV charging loads on the voltage profiles of radial and parallel feeders, resulting in voltage drop across the entire line and a V-curve shape voltage profile. cheng2014evaluating proposes a method for extracting the geographical dispersion of EV integrations and evaluates the reliability at load points. shepero2018modeling focuses on weak points in city-scale by spatio-temporal modeling of PV and EV. Based on voltage distribution and penetration rates, ucer2018learning obtains the conclusion that the congestion condition in the distribution grid can be judged at each local node only by checking its own voltage level. In addition, the issue of attack also deserves consideration; attackers with knowledge of the grid topology can craft smarter attacks by locating the weakest buses that are more likely to cause larger disturbances once attacked, to disrupt the system with smaller numbers of compromised EVs sayed2022electric .
4.2 Real-time management system
The charging and discharging patterns of EVs are largely influenced by the unpredictable travel habits of their users. As EVs are integrated into the distribution system in an unpredictable way, significant uncertainties arise in the timing and location of grid loads. This unpredictability complicates load forecasting and increases the challenges of real-time grid monitoring and management. To address these issues, optimal charging strategies are often designed to maximize the use of renewable energy sources to meet the fluctuating demands of EVs. However, renewable energy generation is intermittent and variable, further aggravating uncertainties in real-time management. To address these issues, many studies have focused on handling the uncertainties of EV charging, renewable energy generation, real-time electricity price, and user behavior, where methods of stochastic optimization kaewdornhan2023real , scenario-based stochastic programming yan2023real ; eddahech2012real ; yao2016real , robust optimization zhao2015risk ; vaya2015self , and dynamic programming mohamed2013real ; you2016real ; binetti2015scalable ; liao2015dispatch have been widely applied. Though these approaches can effectively mitigate the impact of uncertainties, many face the challenge of computational dimensionality. For instance, capturing the characteristics of uncertainty requires numerous scenarios, which leads to a heavy computational burden. Even at a moderate scale, the stochastic programming for EV charging dispatching needs several hours to solve yan2023real . To address the problem of high computational intensity, decentralized and distributed methods have received considerable attention chokkalingam2017real ; csengor2020real ; li2023constrained . Compared to centralized schemes, decentralized approaches provide better scalability and real-time performance li2023constrained . li2023constrained proposes a decentralized framework and a dynamic energy boundary model of individual EVs to achieve discreteness of charging demands, which needs to train only a single agent model to enable local decision-making for multiple charging piles within a charging station. binetti2015scalable utilizes the scalable real-time greedy algorithm, zhou2020admm adopts alternating direction method of multipliers to schedule EVs in a decentralized method, and latifi2018agent designs a game theoretic model to keep both the local optimality and the global optimality for the aggregator. However, the aforementioned distributed methods have limited capability in addressing system-level objectives and overall cooperation requirements. Many studies focus on addressing the balance between mitigating the impact of uncertainties and high computational intensity. wu2017two proposes two control algorithms, SPLET and SAA_SPLET, to reduce the number of scenarios and alleviate the computational burden. li2023constrained introduces a dynamic energy boundary model for EVs to meet charging demand without penalty terms. wang2019reinforcement designes four feature functions to approximate the state-action function, reducing the dimension of the state space. The issue of EV charging and discharging dispatching is essentially a time arrangement for EV charging and discharging control while considering various uncertainties li2023constrained . The current control strategy will influence the system’s next state, which increases computational complexity and affects the convergence speed shen2021real . Therefore, formulating EV charging control issues as a Markov Decision Process (MDP) is a practical approach, where Model-free Deep Reinforcement Learning (DRL) techniques can be used as solutions for sequential decision-making problems zhang2018review to find optimal strategies in complex and uncertain environments.
Due to the randomness of traffic conditions, user behavior, real-time electricity price, the arrival and departure time of EVs, as well as battery states and charging demands, are dynamic and challenging to predict accurately, which makes the management of EV charging challenging wan2018model . In recent years, many day-ahead scheduling methods zhao2015risk ; vaya2015self ; erdinc2014smart have been used to address this issue. However, they have limited capability in handling more complex real-time scenarios involving time-varying EV charging demand and electricity price. Real-time management systems with high efficiency and accuracy, along with appropriately designed DR mechanisms, are two key factors for DSOs (Distribution System Operators) in managing EV charging and discharging. The DR mechanism has already received extensive attention in akhavan2017managing ; ghavami2014nonlinear ; chekired2017decentralized , while there are still mismatches between day-ahead scheduling and real-time demand. Apart from the conventional solution of quickly starting up backup generators to provide a power supply, the EV battery system can also serve as a solution by treating EVs as flexible energy storage units capable of discharging electricity back into the grid. Besides, the EV V2B (Vehicle-to-Building) and SAB (Smart Aggregator-Based) systems can also be used as strategies to reduce operational costs wu2017two . Therefore, many studies have focused on enhancing the accuracy and efficiency of real-time management. For instance, in zhao2015risk , the state-action function is represented by a linear combination of feature functions to transform the decision in the time-varying action space into four time-invariant constants, solving the difficulty in representing the time-varying state and action space in EV charging. jiang2019real formulates the charging optimization problem as a cost minimization problem and proposes an improved binary grey wolf optimizer to improve the convergence speed and optimization accuracy. long2021efficient develops an ordinal optimization-based method to search for optimal charging strategies within seconds while providing a performance guarantee. yang2021dynamic introduces a stochastic dynamic simulation modeling framework for EV fast-charging real-time management, incorporating a multi-server queueing model and a multinomial logit station choice model based on charging prices, expected wait times, and detour distances. csengor2020real uses a linear programming method to propose an energy management model for electric vehicle parking lots based on real-time optimization to maximize load factor. yan2023real develops a real-time feedback integrated online algorithm based on Lyapunov optimization, which provides theoretical bounds for maximum charging delays.
Figure 5: Framework of real-time management system of EV charging.
As large-scale EVs integrate into the power grid and the continuous development of real-time energy management systems, the scope of EV charging has expanded significantly. The coordination of bi-directional energy flows between EV charging systems and other systems, such as photovoltaic, thermal energy, and energy storage systems, has gained widespread attention. Many studies focus on real-time energy management, integrating EV charging with multiple interconnected systems. zhang2018dynamic proposes a dynamic cost optimization scheduling method based on real-time information of EV charging demand and PV generation to control the charging process of each EV in a parking lot without considering ESS. shen2021real introduced a Combined Heat and Power (CHP) system and proposed a real-time energy management algorithm for a microgrid based on the Lyapunov optimization technique to minimize the average cost of the microgrid. jiang2019real investigates the charging optimization for EVs in parking lots integrated with ESS and PV systems and proposes a real-time EV charging dispatching strategy based on an improved binary grey wolf optimizer. Figure 5 presents the framework of a real-time management system of EV charging. The Summary of the studies on the Real-time management system of EV charging is presented in Table 2.
Table 2: Summary of the studies on real-time management system of EV charging.
Reference Year Technical foundation Implementation approach Application
yang2021dynamic 2021 Multi-server queueing model and EV user behavior model Develop a stochastic dynamic simulation modeling framework of EV fast-charging stations Real-time management and strategic planning
wan2018model 2019 Model-free approach based on deep reinforcement learning Extract discriminative features from the electricity prices and approximate the optimal action-value function through a Q network. Optimal strategy for real-time EV charging scheduling
jiang2019real 2019 Improved binary grey wolf optimizer Short-term PV power prediction and IBGWO Real-time EV charging scheduling strategy
long2021efficient 2021 Aggregated PEV charging model based on the incomplete Beta function Parameterize the aggregated charging policy using the energy boundaries to express the charging flexibility Real-Time EV charging scheduling
shen2021real 2021 Lyapunov stochastic optimization technique Optimize the microgrid average cost without knowing future price, demands, and other system information Real-time energy management algorithm for EV-based microgrid
wang2019reinforcement 2021 Reinforcement learning Develop a model-free data-driven method for EV charging stations with random EV arrivals and departures Joint pricing and charging scheduling
csengor2020real 2021 Linear programming Propose a peak load limitation oriented DR program with the objective of maximizing load factor Real-time energy management model for EV parking lot
li2023constrained 2023 Model-free DRL Propose a decentralized framework, a dynamic energy boundary model and formulate a Markov Decision Process Large-scale real-time EV scheduling
yu2018real 2018 Convex optimization problem of energy scheduling Maximize consumption of new energy generation considering the inhomogeneous EV load and real-time price market Real-time energy management for EV charging station with RE and ESS
wan2018data 2018 Data-driven approach based on deep reinforcement learning Formulate EV charging management as a Markov Decision Process which has unknown transition probability Optimal EV charging strategy
manbachi2016impact 2016 Real-time simulator and monitoring platform through DNP.3 protocol Introduce an AMI-based VVO engine to minimize grid loss and Volt–VAR control assets costs while maximizing CVR benefit Real-time co-simulation platform
kaewdornhan2023real 2023 Multi-agent deep reinforcement learning optimization Tackle the multi-home energy management problem with EV charging and discharging scheduling considering uncertainties Real-time multi-home energy management with EV charging scheduling
yan2023real 2024 Lyapunov optimization method Propose an offline model with bounds of the aggregate EV power flexibility region, an online algorithm with a theoretical bound for the maximum charging delay and real-time feedback design Real-time feedback based online aggregate EV power flexibility characterization
yao2016real 2017 Combination of linear programming and modified convex relaxation Coordinate EV charging loads and accommodate DR programs. Real-time EV charging scheme in the parking station
binetti2015scalable 2015 Scalable real-time greedy algorithm Schedule a large population of EVs in a decentralized fashion, considering the discrete charging scenario Large-scale EVs scheduling
liao2015dispatch 2015 Stochastic dynamic programming Determine optimal dispatch schedule with different temporal variations of energy production and consumption Economic dispatch of the controllable resources in an EV charging station
zhou2020admm 2021 Alternating direction method of multipliers Model the EV fast charging problem as an optimization coordination problem subject to coupled feeder capacity constraints in the distribution network Optimal strategy profile for EVs
latifi2018agent 2019 Decentralized game theoretic model Minimize customers’ payments, maximize grid efficiency, and provide the maximum potential capacity for ancillary services Real-time dynamic pricing model
5 Conclusion and future outlook
This paper first discusses the impact of EV charging on the operation of power distribution systems from the perspective of power quality and grid stability. For the issue of real-time stable operation, the coping capacity of EV charging areas to grid faults and solutions to peak loads are reviewed. Further in-depth analysis is carried out to explore how to reduce the negative impact of EV charging on the distribution network and guide positive impacts, followed by a comprehensive summary of specific treatment methods. Considering that the state of charge of EV batteries and charging demand are difficult to predict, widespread random integration of EVs will introduce significant uncertainties into the real-time operation of the grid, increasing the complexity of real-time management. Therefore, this paper conducts an in-depth analysis of the real-time management of EV charging from the perspectives of monitoring and dispatching. First, real-time state estimation based on the topology of the distribution network is performed to assess the sensitivity of different locations in the grid to accommodate EV loads. Further strengthening of monitoring and management of weaknesses in the distribution network during operation is also necessary. Given the increasingly complex real-time scenarios of EV charging demand and price fluctuations, day-ahead scheduling methods have limited capabilities. While real-time scheduling methods usually face significant computational burdens when addressing uncertainties, balancing the impacts of uncertainties and computational intensity is crucial. Decentralized methods are effective for solving computationally intensive problems. However, they have limited capacity to address system-level goals and cooperative demands, leading to widespread attention to model-free deep reinforcement learning technologies. Moreover, it’s important to study highly efficient and accurate real-time management systems through improved algorithms, mechanisms, and connections with other systems. Regarding future research directions, enhancing the coping capacity of EV charging areas to grid faults is worthy of more attention. Additionally, the balance between the impacts of uncertainties and computational intensity needs to be further explored to improve the accuracy and efficiency of EV real-time management.
[…]